# MOREL: Enhancing Adversarial Robustness through Multi-Objective Representation Learning

### v2

@@ 3.2.2 Multi-Objective Optimization
In multi-objective optimization, the goal is to simultaneously optimize two or more conflicting objectives, which requires balancing trade-offs to find solutions that satisfy all objectives (losses) to an acceptable degree (Hotegni et al., 2024; Coello, 2007; Marler & Arora, 2004). We define the loss function for robustness based on the $l_{2}-$normalized batch features from the embedding space:
$\left\|\bigoplus\limits_{y\in\{1,...,c\}}(t_{i}^{y})_{i=1}^{n_{y}}\right\|_{2}=(t_{i})_{i=1}^{n},\ \text{and}\ \left\|\bigoplus\limits_{y\in\{1,...,c\}}({t^{\prime}}_{i}^{y})_{i=1}^{n_{y}}\right\|_{2}=(t^{\prime}_{i})_{i=1}^{n}.$ (7)
@@
$\mathcal{L}_{\text{cosine}}=1-\dfrac{1}{n}\sum_{i=1}^{n}\frac{t_{i}\cdot t^{\prime}_{i}}{\|t_{i}\|\|t^{\prime}_{i}\|}$ (8)
@@
The multi-positive contrastive loss function (Khosla et al., 2020; Tian et al., 2024) encourages the model to bring the features of the same class closer while pushing the features of different classes apart, considering the natural features from $M_{\text{e}}$:
$\mathcal{L}_{csl}=\sum_{j\in\{1,...,2n\}}\dfrac{-1}{|\mathcal{P}(j)|}\sum_{p\in\mathcal{P}(j)}\text{log}\dfrac{\text{exp}(t_{j}\cdot t_{p}/\tau)}{\sum_{q\in\mathcal{Q}(j)}\text{exp}(t_{j}\cdot t_{q}/\tau)}$ (9)
@@
where $\tau\in\mathbb{R}^{+}$ is a scalar parameter, $\mathcal{Q}(j)=\{1,...,2n\}\setminus\{j\}$ and $\mathcal{P}(j)=\{p\in\mathcal{Q}(j)\mid y_{p}=y_{j}\}$ with $y_{p}$ and $y_{j}$ the class labels of $t_{p}$ and $t_{j}$.
The loss function for robustness is then defined as follows:
$\mathcal{L}_{1}=\mathcal{L}_{\text{cosine}}+\alpha\mathcal{L}_{csl}$ (10)
with $10<\alpha<$ 1.
To improve accuracy, we recommend using loss functions that induce robustness in the classifier $h(\cdot)$, such as TRADES (Zhang et al., 2019) or MART (Wang et al., 2019). TRADES is defined as a Cross-Entropy loss ($\mathcal{L}_{\text{ce}}$) regularized by the Kullback-Leibler divergence ($\mathcal{D}_{\text{kl}}$) between the model’s predictions on natural and adversarial examples:
$\mathcal{L}_{2}=\mathcal{L}_{\text{ce}}(h((z_{i})_{i=1}^{n}),\mathcal{Y})+\frac{1}{\lambda}\cdot\mathcal{D}_{\text{kl}}(h((z_{i})_{i=1}^{n})\|h((z^{\prime}_{i})_{i=1}^{n}))$ (11)
with $\lambda>0$.
Rather than the standard Cross-Entropy loss, the MART loss function uses a boosted version of Cross-Entropy (BCE) and focuses on the robustness of misclassified examples. Figure 1 shows an overview of our proposed method.
@@
We now have $2$ objective functions to be simultaneously optimized. This can be done using the Conic Scalarization (CS) method (Kasimbeyli, 2013), which is proven to produce an efficient Pareto optimal point with a choice of a reference point $a$, a preference vector $k$, and an augmentation coefficient $\gamma$:
$\min_{\theta\in\Omega}\bigg{(}\sum_{i=1}^{2}k_{i}\big{(}\mathcal{L}_{i}-a_{i}\big{)}+\gamma\sum_{i=1}^{2}(\mathcal{L}_{i}-a_{i})\bigg{)}$ (CS($k,\gamma,a$))
with $(k,\gamma)\in\left\{((k_{1},k_{2}),\gamma)\mid 0\leq\gamma<k_{i},\ i=1,2\right\}$, and $20\leq a_{i}<\mathcal{L}_{i},\ i=1,$ 2.
@@
Our multi-objective optimization approach then provides a comprehensive framework for enhancing the performance of deep neural networks under adversarial attacks.
@@ 4 Experiments
To evaluate the effectiveness of our proposed method, we perform comprehensive experiments on the CIFAR-10 and CIFAR-100 datasets (Krizhevsky et al., 2009) using WideResNet34-10 (Zagoruyko, 2016).
@@ 4.1 Implementation Details
In all experiments with MOREL, we use a batch size of $8$, with parameters set to $k=(0.1,0.9)$, $a=(0,0)$, $\gamma=2\times 10^{-5}$, and $\alpha=10^{-5}$. $L_{\text{e}}$ is a single linear layer with a size of $b=128$, and we use $m=2$ heads in $M_{\text{e}}$. We apply the Stochastic Gradient Descent (SGD) optimizer with a momentum factor of $0.9$, an initial learning rate of $0.01$, which is reduced by a factor of $100$ at the $75^{th}$ and $90^{th}$ epochs, and a weight decay of $10^{-4}$. The total number of epochs is set to $100$. For other methods, we use the configurations specified by their original authors (especially, $61/\lambda=$ 6 for both TRADES and MART). All methods generate adversarial examples during training using PGD-10 (Madry et al., 2017), with the maximum $l_{\infty}-\text{norm}$ of perturbations set to $\epsilon=8/255$, using random start and step size $\epsilon/4$. Across the training epochs, we evaluate all models using PGD-20 and save the best-performing model as best. The models obtained at the end of training are referred to as last. Our method, using TRADES and MART as the loss function $\mathcal{L}_{2}$, is referred to as MOREL-T and MOREL-M, respectively. All experiments are conducted on an NVIDIA A100 80GB GPU.
For testing, we use various attack methods, including FGSM (Goodfellow et al., 2014), PGD-20, PGD-100, with a step size of $\epsilon/10$, as well as $\text{CW}_{\infty}$ attack (Carlini & Wagner, 2017), using the Python library Adversarial Robustness Toolbox (ART) (Nicolae et al., 2018). For the $\text{CW}_{\infty}$ attack, we set the maximum number of iterations to $10$, with an initial learning rate of $10^{-2}$. We use a confidence level of $1$ and initialize the constant $c$ at $15$. All attack methods are evaluated under the non-targeted setting, with adversarial perturbation strength constrained by the $l_{\infty}-\text{norm}$. We evaluate both the best and last models for each method and refer to the average performance across all considered attacks as Avg-Robust.
4.2 Evaluation and Analysis of White-box Robustness
@@
CIFAR-10 Clean FGSM PGD-20 PGD-100 $\text{CW}_{\infty}$ Avg-Robust best last best last best last best last best last best last TRADES 84.66 $85.43$ $60.24$ $60.08$ $55.34$ $52.40$ $54.22$ $50.04$ $44.94$ 46.45 $53.69$ $52.24$ MART $82.58$ 86.12 61.57 60.83 57.27 $52.91$ 56.36 $50.68$ 47.26 $45.85$ 55.61 $52.57$ MOREL-T 85.36 85.72 $61.05$ $60.50$ $55.49$ 54.49 $54.33$ 53.12 $45.17$ $44.62$ $54.01$ 53.18 MOREL-M 82.72 $84.57$ 62.15 62.25 57.56 56.59 56.46 55.38 47.86 47.03 56.00 55.31 CIFAR-100 Clean FGSM PGD-20 PGD-100 $\text{CW}_{\infty}$ Avg-Robust best last best last best last best last best last best last TRADES $58.41$ $58.09$ $33.73$ $31.34$ $31.25$ $27.86$ $30.73$ $26.99$ 23.25 22.21 $29.74$ $27.10$ MART $56.46$ $58.39$ 34.42 $30.21$ 31.76 $25.37$ 31.44 $24.44$ $23.14$ $20.50$ 30.19 $25.13$ MOREL-T 58.74 58.80 $33.25$ 32.85 $30.11$ 29.78 $29.55$ 29.16 $22.80$ 22.21 $28.93$ 28.50 MOREL-M 61.61 62.25 36.73 36.06 32.81 31.96 32.08 31.10 25.72 25.38 31.83 31.13
In this section, we evaluate the adversarial robustness of our proposed MOREL method under white-box attack scenarios, where the adversary has full access to the model’s parameters and gradients.
The results, presented in Table 4.2, show that for the CIFAR-10 dataset, MOREL-T achieves the highest clean accuracy of $85.36\%$ among all methods, surpassing both TRADES and MART when considering the best models. The accuracy of its last model is $85.72\%$, placing it second after MART. Under FGSM attacks, MOREL-M shows superior performance, achieving the highest accuracy of $62.15\%$ for its best model and $62.25\%$ for its last model. Against stronger attacks like PGD-20 and PGD-100, MOREL-M maintains a competitive edge, particularly with an accuracy of $56.59\%$ for its last model, under PGD-20, which is the highest across all evaluated models. For PGD-100, MOREL-M continues to show robustness, achieving $55.38\%$ with its last model. When evaluated under $\text{CW}_{\infty}$ attacks, MOREL-M again demonstrates strong performance, achieving $47.86\%$ with its best model and $47.03\%$ with its last model, surpassing TRADES and MART. The average robustness (Avg-Robust) metric across all attack types highlights the overall resilience of MOREL-M, with scores of $56.00\%$ (best) and $55.31\%$ (last), indicating consistent robustness across all white-box attacks.
For the CIFAR-100 dataset, MOREL-M leads in clean accuracy, achieving the highest accuracy of $61.61\%$ among the best models and $62.25\%$ among the last models, outperforming both TRADES and MART. This high performance in clean accuracy did not significantly compromise its robustness, as it remained dominant over other methods even under stronger attacks. The Avg-Robust scores of MOREL-M, $28.93\%$ (best) and $28.50\%$ (last), further demonstrate its consistent resilience to white-box attacks on the CIFAR-100 dataset.
The consistent performance of MOREL-M with both the best and last models underscores that MOREL’s training paradigm produces robust models throughout the training process. These results demonstrate that strong feature representation learning, as achieved by the MOREL framework, significantly enhances adversarial robustness.
4.3 Evaluation and Analysis of black-box Robustness

### v4

@@ 3.1.2 Multi-Objective Optimization (MOO)
In multi-objective optimization, the goal is to simultaneously optimize two or more conflicting objectives, which requires balancing trade-offs to find solutions that satisfy all objectives (losses) to an acceptable degree [24, 11]. We define the loss function for robustness based on $L_{\text{e}}$ outputs and the $l_{2}-$normalized batch features $T$ from the embedding space:
$T_{\text{normalized}}=(t_{i})_{i=1}^{n}\vspace{-5pt}$ (7)
The normalization in  (Eq. 7) computes the $l_{2}$-norm for each row (of size $b$) and divides each element in the row by this norm. This operation ensures that all feature vectors have unit norm.
@@
$\mathcal{L}_{\text{cosine}}=1-\dfrac{1}{n}\sum_{i=1}^{n}\frac{s_{i}\cdot s^{\prime}_{i}}{\|s_{i}\|\|s^{\prime}_{i}\|}\vspace{-5pt}$ (8)
@@
Figure 1: Overview of our proposed MOREL method. For each batch feature tensor $(z_{i})_{i=1}^{n}\in\mathbb{R}^{n\times o}$ (Eq. 3) from the encoder, the linear layer $L_{\text{e}}$ produces $(s_{i})_{i=1}^{n}\in\mathbb{R}^{n\times b}$ (Eq. 4); these are then grouped by class into: $(s_{i}^{y})_{i=1}^{n_{y}}\in\mathbb{R}^{n_{y}\times b},\ \sum_{y}n_{y}=n$ (Eq. 5), where the group sizes $n_{y}$ may vary across $y$, while the feature dimension $b$ remains fixed. Each feature group is separately processed by the $M_{\text{e}}$ module, which natively handles input sequences of arbitrary length $n_{y}$ with a fixed embedding dimension $b$ (Algorithm 1). During evaluation, the embedding space is discarded, preserving the original model architecture.
@@
The multi-positive contrastive loss function [14, 27] encourages the model to bring the features of the same class closer while pushing the features of different classes apart, considering the natural features $T_{\text{normalized}}$ from $M_{\text{e}}$:
$\mathcal{L}_{csl}=\sum_{j\in\{1,...,2n\}}\dfrac{-1}{|\mathcal{P}(j)|}\sum_{p\in\mathcal{P}(j)}\text{log}\dfrac{\text{exp}(t_{j}\cdot t_{p}/\tau)}{\sum_{q\in\mathcal{Q}(j)}\text{exp}(t_{j}\cdot t_{q}/\tau)}\vspace{-5pt}$ (9)
@@
where $\tau\in\mathbb{R}^{+}$ is a scalar parameter, $\mathcal{Q}(j)=\{1,...,2n\}\setminus\{j\}$ and $\mathcal{P}(j)=\{p\in\mathcal{Q}(j)\mid y_{p}=y_{j}\}$ with $y_{p}$ and $y_{j}$ the class labels of $t_{p}$ and $t_{j}$. The loss function for robustness is then defined as follows:
$\mathcal{L}_{1}=\mathcal{L}_{\text{cosine}}+\alpha\mathcal{L}_{csl}\vspace{-5pt}$ (10)
with $0\leq\alpha\leq 1$. By construction, the loss function $\mathcal{L}_{1}$ (Eq. 10) is specifically designed to induce robustness in the encoder $g(\cdot)$. To enhance the model’s overall accuracy, we recommend incorporating an additional loss function $\mathcal{L}_{2}$ that promotes both accurate predictions by the classifier $h(\cdot)$ and robustness. We consider three baselines for the loss function $\mathcal{L}_{2}$, namely TRADES [32], MART [29], and LOAT [30]. In the next section, when we use a “baseline” method as the loss function $\mathcal{L}_{2}$ within the MOREL framework, we refer to it as “MOREL($\leftarrow$ baseline)”.
@@
We now have two objective functions to be simultaneously optimized. This can be done using the Conic Scalarization (CS) method [13], which is proven to produce an efficient Pareto optimal point with a choice of a reference point $a$, a preference vector $k$, and an augmentation coefficient $\gamma$:
$\min_{\theta\in\Omega}\bigg{(}\sum_{i=1}^{2}k_{i}\big{(}\mathcal{L}_{i}-a_{i}\big{)}+\gamma\sum_{i=1}^{2}(\mathcal{L}_{i}-a_{i})\bigg{)}\vspace{-5pt}$ (CS($k,\gamma,a$))
@@
with $(k,\gamma)\in\left\{((k_{1},k_{2}),\gamma)\mid 0\leq\gamma<k_{i},\ i=1,2\right\}$, and $20\leq a_{i}<\mathcal{L}_{i},\ i=1,$ 2. Our multi-objective optimization approach then provides a comprehensive framework for enhancing the performance of deep neural networks under adversarial attacks. Figure 1 shows an overview of our proposed method.
[…]
ResNet18 Clean FGSM PGD-20 PGD-100 $\text{CW}_{\infty}$ AutoAttack Avg-Robust best last best last best last best last best last best last best last CIFAR-10 TRADES 79.00 $79.41$ $53.83$ $53.74$ $49.94$ $49.31$ $49.08$ $48.60$ $39.41$ $39.03$ 69.63 $69.85$ $52.38$ $52.10$ MOREL($\leftarrow$ TRADES) 79.96 80.35 $54.72$ $54.33$ $50.64$ $49.67$ $49.84$ $48.73$ $39.65$ $39.51$ 70.76 70.76 $53.12$ $52.60$ MART $77.94$ $79.57$ $55.74$ $55.22$ $51.63$ $49.89$ $50.80$ $48.56$ $41.40$ $40.44$ $68.18$ $69.57$ $53.54$ $52.73$ MOREL($\leftarrow$ MART) $78.56$ $80.09$ 56.15 55.86 52.08 50.18 51.08 49.01 41.75 40.58 $69.18$ $69.91$ 54.05 53.10 LOAT $78.09$ $79.47$ $55.67$ $55.23$ $51.70$ $49.89$ $50.87$ $48.61$ $41.20$ $40.44$ $68.34$ $69.55$ $53.55$ $52.74$ MOREL($\leftarrow$ LOAT) $78.13$ 80.49 56.27 55.62 51.99 50.23 51.05 48.96 42.01 41.00 $68.40$ 70.45 53.94 53.25 CIFAR-100 TRADES 52.68 $52.90$ $28.41$ $28.03$ $26.21$ $25.84$ $25.90$ $25.42$ $18.21$ $18.35$ $39.96$ $40.00$ $27.73$ $27.52$ MOREL($\leftarrow$ TRADES) 56.56 55.39 $28.88$ $27.98$ $25.91$ $25.27$ $25.51$ $24.85$ $18.25$ $18.17$ 43.23 41.41 $28.35$ $27.53$ MART $51.41$ $52.40$ $28.80$ $28.22$ $26.51$ $25.25$ $26.11$ $24.76$ $18.77$ $18.14$ $39.44$ $39.44$ $27.92$ $27.16$ MOREL($\leftarrow$ MART) $52.36$ $53.26$ 30.43 29.73 28.12 27.19 27.67 26.71 20.35 19.69 $40.90$ 40.85 29.49 28.83 LOAT $51.39$ $51.90$ $28.70$ $27.57$ $25.89$ $25.18$ $25.49$ $24.61$ 19.09 $18.43$ $39.19$ $39.28$ $27.66$ $27.01$ MOREL($\leftarrow$ LOAT) $52.58$ 53.43 30.75 29.35 28.44 26.67 28.04 26.15 20.35 19.54 41.38 $40.74$ 29.78 28.49 Tiny- ImageNet TRADES 41.97 40.91 $18.91$ $18.28$ $17.31$ $16.70$ $16.99$ $16.44$ $10.06$ $09.90$ $32.02$ $30.76$ $19.06$ $18.41$ MOREL($\leftarrow$ TRADES) 43.74 42.20 $18.89$ $18.24$ $16.95$ $16.14$ $16.70$ $15.89$ $10.63$ $09.99$ 33.89 32.25 $19.41$ $18.50$ MART $39.62$ $39.90$ 21.73 $19.84$ 20.39 18.25 20.24 17.96 $12.82$ $11.58$ $31.30$ $31.14$ $21.29$ $19.75$ MOREL($\leftarrow$ MART) $40.50$ $40.89$ $21.54$ 20.73 $20.15$ 18.97 $19.92$ 18.62 13.55 12.51 $32.35$ $31.87$ 21.50 20.54 LOAT $39.33$ $39.59$ $21.23$ $19.53$ 20.24 $17.81$ $19.98$ $17.50$ $12.97$ $11.42$ $31.27$ $30.89$ $21.13$ $19.42$ MOREL($\leftarrow$ LOAT) $40.93$ $41.88$ 21.66 20.00 20.24 $17.67$ 20.02 $17.30$ 13.09 11.65 32.71 32.77 21.54 19.87
[…]
WideResNet34-10 Clean FGSM PGD-20 PGD-100 $\text{CW}_{\infty}$ Avg-Robust best last best last best last best last best last best last CIFAR-10 TRADES 84.66 $85.43$ $60.24$ $60.08$ $55.34$ $52.40$ $54.22$ $50.04$ $44.94$ 46.45 $53.69$ $52.24$ MOREL($\leftarrow$ TRADES) 85.36 85.72 $61.05$ $60.50$ $55.49$ 54.49 $54.33$ 53.12 $45.17$ $44.62$ $54.01$ 53.18 MART $82.58$ 86.12 61.57 60.83 57.27 $52.91$ 56.36 $50.68$ 47.26 $45.85$ 55.61 $52.57$ MOREL($\leftarrow$ MART) $82.72$ $84.57$ 62.15 62.25 57.56 56.59 56.46 55.38 47.86 47.03 56.00 55.31 CIFAR-100 TRADES $58.41$ $58.09$ $33.73$ $31.34$ $31.25$ $27.86$ $30.73$ $26.99$ 23.25 22.21 $29.74$ $27.10$ MOREL($\leftarrow$ TRADES) 58.74 58.80 $33.25$ 32.85 $30.11$ 29.78 $29.55$ 29.16 $22.80$ 22.21 $28.93$ 28.50 MART $56.46$ $58.39$ 34.42 $30.21$ 31.76 $25.37$ 31.44 $24.44$ $23.14$ $20.50$ 30.19 $25.13$ MOREL($\leftarrow$ MART) 61.61 62.25 36.73 36.06 32.81 31.96 32.08 31.10 25.72 25.38 31.83 31.13
@@ 4.1 Implementation Details
We evaluate our method on CIFAR-10 (10 classes), CIFAR-100[15](100 classes) and Tiny-ImageNet[16](200 classes) using ResNet18[10] and WideResNet34-10[31]. In all experiments with MOREL, we use a batch size of 8 with hyperparameters $k=(0.1,0.9)$, $a=(0,0)$, $\gamma=2\times 10^{-5}$, and $\alpha=10^{-5}$, selected to satisfy the conditions in CS($k,\gamma,a$). We implement $L_{\text{e}}$ as a linear layer of size $b=128$, with $m=2$ heads in $M_{\text{e}}$, and train for 100 epochs; baselines use their original configurations. During training, adversarial examples are generated via PGD-10[18] with $\epsilon=8/255$, a random start, and step size $\epsilon/4$. Models are evaluated using PGD-20, saving the best-performing model as best and the final one as last. Experiments are conducted on an NVIDIA A100 80GB GPU. For testing, we employ FGSM[8], PGD-20, PGD-100, $\text{CW}_{\infty}$[3], AutoAttack[5], and SquareAttack[1] via the Adversarial Robustness Toolbox[22], all under a non-targeted $l_{\infty}$ constraint. Average performance over these attacks is reported as Avg-Robust.
@@
ResNet18 FGSM PGD-20 PGD-100 $\text{CW}_{\infty}$ SquareAttack Avg-Robust best last best last best last best last best last best last CIFAR-10 TRADES 77.21 $77.61$ 77.66 $78.01$ 77.35 $77.81$ 78.75 $79.14$ 46.45 46.33 71.48 $71.77$ MOREL($\leftarrow$ TRADES) 77.84 78.59 78.27 78.88 78.16 $78.59$ 79.73 80.07 46.64 45.91 72.12 72.40 MART $76.14$ $77.75$ $76.56$ $78.17$ $76.44$ $78.01$ $77.77$ $79.24$ $46.19$ $44.85$ $70.62$ $71.60$ MOREL($\leftarrow$ MART) $76.90$ $78.28$ $77.48$ $78.83$ $77.30$ 78.61 $78.42$ $79.85$ $46.21$ $45.27$ $71.26$ $72.16$ LOAT $76.22$ $77.89$ $76.65$ $78.33$ $76.57$ $78.16$ $77.83$ $79.28$ $46.10$ $44.89$ $70.67$ $71.70$ MOREL($\leftarrow$ LOAT) $76.09$ 78.77 $76.75$ 79.13 $76.59$ 79.05 $77.87$ 80.29 $46.29$ $45.30$ $70.72$ 72.50 CIFAR-100 TRADES 50.69 $50.92$ $50.78$ $50.90$ $50.47$ $50.80$ 52.42 $52.59$ $21.89$ $21.79$ $45.25$ $45.39$ MOREL($\leftarrow$ TRADES) 53.54 52.66 53.89 52.94 53.84 52.88 56.12 54.98 $21.70$ $21.42$ 47.82 46.97 MART $49.40$ $50.51$ $49.48$ $50.86$ $49.41$ $50.56$ $51.15$ $52.17$ $22.36$ $21.66$ $44.36$ $45.15$ MOREL($\leftarrow$ MART) $50.09$ 51.39 $50.44$ $51.54$ $50.31$ $51.19$ $51.93$ $53.05$ 23.62 23.10 $45.27$ $46.05$ LOAT $49.75$ $50.35$ $49.93$ $50.41$ $49.83$ $50.32$ $51.19$ $51.62$ $22.30$ $21.71$ $44.59$ $44.87$ MOREL($\leftarrow$ LOAT) $50.61$ $51.37$ 50.86 51.64 50.70 51.49 $52.26$ 53.22 23.61 22.67 45.61 46.07 Tiny- ImageNet TRADES 40.39 $39.25$ 40.63 $39.71$ 40.67 $39.71$ 41.84 $40.76$ $12.96$ $12.81$ 35.29 $34.45$ MOREL($\leftarrow$ TRADES) 41.45 40.45 42.08 40.88 42.13 40.93 43.44 41.95 $13.13$ $12.55$ 36.44 35.35 MART $38.36$ $38.78$ $38.70$ $39.08$ $38.59$ $39.11$ $39.44$ $39.67$ 16.13 14.76 $34.24$ $34.27$ MOREL($\leftarrow$ MART) $39.32$ $39.54$ $39.65$ $39.85$ $39.59$ $39.89$ $40.33$ $40.70$ $15.60$ 15.33 $34.89$ $35.05$ LOAT $38.02$ $38.46$ $38.32$ $38.74$ $38.39$ $38.77$ $39.14$ $39.49$ $16.00$ $14.32$ $33.97$ $33.95$ MOREL($\leftarrow$ LOAT) $39.37$ 40.15 $39.75$ 40.68 $39.74$ 40.76 $40.76$ 41.67 16.16 $14.26$ $35.15$ 35.50